OpenStudy (anonymous):

WILL GIVE MEDAL AN FAN :) How many neutrons are in Cesium-130 (Cs)? 55 75 130 185

OpenStudy (anonymous):

@HazelLuv99 Do you know how to do this?

OpenStudy (anonymous):

@samii22 @sarahnichols3

OpenStudy (hazelluv99):

i'm not sure srry

OpenStudy (irishboy123):

Caesium 135 has 135 - 55 = 80 neutrons.

OpenStudy (anonymous):

but thats not an answer choice

OpenStudy (anonymous):

oh wait sorry is it 55.. lol my brain is still on sleep mode......

OpenStudy (anonymous):

How many protons are in Cesium-135 (Cs)? 35 55 80 195 heres another one i dont understand these @IrishBoy123

OpenStudy (anonymous):

are they both 55

OpenStudy (anonymous):

@Tootles143

OpenStudy (anonymous):

Woops. Sowy hun. I am not at all good at this. :'(

OpenStudy (anonymous):

okay thank you though:)

OpenStudy (irishboy123):

SORRY: lets be clear: How many neutrons are in Caesium-130 (Cs)? Caesium 130 has 130 - 55 = 75 neutrons. How many neutrons are in Caesium-130 (Cs)? Caesium 135 has 135 - 55 = 80 neutrons.